OpenStudy (umlika):

prove that x^4/x^2+2 is a parabola.

geerky42 (geerky42):

\[\dfrac{x^4}{x^2}+2\] ?

OpenStudy (umlika):

its x^4 and the numerator is x^2+2 as a whole

OpenStudy (umlika):

no i mean denomenator*

OpenStudy (paxpolaris):

\[{x^4 \over x^2+2}\] so like this ^^^ ?

OpenStudy (umlika):

yea :)

geerky42 (geerky42):

Well, \(\dfrac{x^4}{x^2+2}\) is not polynomial function.

OpenStudy (umlika):

yeah but my teacher said that it is a parabola and we have to prove it why

OpenStudy (paxpolaris):

what is the definition of a parabola?

geerky42 (geerky42):

If it is not polynomial function, then it cannot be parabola. Are you sure it's not \(\dfrac{x^4}{x^2}+2\)?

geerky42 (geerky42):

I guess definition of parabola is not really what I think it is

OpenStudy (umlika):

umm... i m preety sure that the question is what i told you, but yeah this is confusing, when i plot this on a graphing calculator it showed parabola

geerky42 (geerky42):

It looks like parabola, but it is not parabola, according "normal" definition of parabola. Are you given definition? What is definition of parabola?

OpenStudy (umlika):

No, just this :/

OpenStudy (anonymous):

According to http://mathworld.wolfram.com/Parabola.html a parabola is "the set of all points in the plane equidistant from a given line and a given point F not on the line"

OpenStudy (umlika):

so this problem is not a parabola ?

OpenStudy (paxpolaris):

no i don't think so. .. it is very close to a parabola for large x. but not exactly one.